OpenStudy (anonymous):

-12x − 3 =[ ]x − 2 Solution: x = -3

OpenStudy (anonymous):

Uhh, are we supposed to figure out what goes in the brackets??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Whaat?? Try: -12(-3) - 3= __x- 2 36-3 + 2= ___x 35= x ???

OpenStudy (anonymous):

sorry ii cant help

OpenStudy (anonymous):

Its ok

OpenStudy (anonymous):

suppose there is a in the bracket then -12x-3=ax-2a 2a-3=x(a+12) then put 3 instead of x 2a-3=3a+36 then a=-39

OpenStudy (anonymous):

thank you very much